OpenStudy (anonymous):

Page 578# 72 Every student in your history class is requite to present a project in front of the class . Each day , 4 students make their presentations in an order chosen at random by the teacher . You make your presentation on the first day. a ) What is the probability that you are chosen to be the first or second presenter on the first day? b) What is the probability that you are chosen to be second or third one the first day?. Compare your answer with that in part ( a )

OpenStudy (anonymous):

Thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

2 people have 2 different answer, I will post again to see what answer from difference people

OpenStudy (anonymous):

14/24

OpenStudy (supersmart1001):

whats ur question

OpenStudy (anonymous):

the answer not match

OpenStudy (supersmart1001):

I looked at the top and it lookes like kropot helped u

OpenStudy (anonymous):

need explain and part b answer too

OpenStudy (anonymous):

I did but don't have result

OpenStudy (anonymous):

the formula is \[\frac{ n! }{ (n-r )!}\]

OpenStudy (supersmart1001):

openstudy.com/updates/572ee741e4b06f843e1e4f15 XD i found ur question instead

OpenStudy (supersmart1001):

I WILL BE BACK

OpenStudy (anonymous):

some one answer 4! = 24 way select first and second is 3x2 =6 second and third the same 3x2 = 6 than 12 out of 12/24 = 1/2 ?

OpenStudy (anonymous):

I hat to do probability

OpenStudy (anonymous):

7/12 and 1/2 which one ?

OpenStudy (phi):

I don't understand dropout's reasoning.

OpenStudy (anonymous):

I don't know either , I try post more to see another answer

OpenStudy (phi):

Here are all the permutations of 4 numbers. Pretend your name is 1 count how many have 1 in the first or 2nd position and divide by the total number >> perms([1 2 3 4]) ans = 1 2 3 4 2 1 3 4 1 3 2 4 2 3 1 4 3 1 2 4 3 2 1 4 1 2 4 3 2 1 4 3 1 3 4 2 2 3 4 1 3 1 4 2 3 2 4 1 1 4 2 3 2 4 1 3 1 4 3 2 2 4 3 1 3 4 1 2 3 4 2 1 4 1 2 3 4 2 1 3 4 1 3 2 4 2 3 1 4 3 1 2 4 3 2 1

OpenStudy (anonymous):

4 students = 4! = 24 ways than find probability first + probability second I'm not sure how to find probability os first and second

OpenStudy (phi):

the brute force way is to count the number of permutations that has 1 in the first or second column. there are 12 of them 12/24 = ½

OpenStudy (anonymous):

first or second is mean first + second

OpenStudy (anonymous):

you said first is 1/2 ?

OpenStudy (anonymous):

second the same 1/2 ?

OpenStudy (phi):

6 have a 1 in the first column 6 have a 1 in the 2nd column so 12 have a 1 in the first or 2nd column (meaning picked to be first or second)

OpenStudy (anonymous):

a) b) the same answer 1/2?

OpenStudy (phi):

Here are the 12 permutations 1 2 3 4 2 1 3 4 1 3 2 4 3 1 2 4 1 2 4 3 2 1 4 3 1 3 4 2 3 1 4 2 1 4 2 3 1 4 3 2 4 1 2 3 4 1 3 2 with 1 in the first or 2nd column

OpenStudy (phi):

yes 2nd or 3rd is also `½

OpenStudy (anonymous):

ok, I will let you know after my home work correct from teacher , can you help me more?

OpenStudy (phi):

what is the question ?

OpenStudy (anonymous):

let me attach file

OpenStudy (anonymous):

I need help number 74, 75

OpenStudy (anonymous):

74) a= 3 C 2 ? b+ 3P2 ?

OpenStudy (phi):

74 a) would be 3 choose 3 (there are 3 colors, and we choose 3 of them 3C3 is 1 in other words, if we ignore order (we do in a combination), we get a red, green, blue

OpenStudy (anonymous):

\[a) \frac{ 3! }{(3-2)!\times2! }\]

OpenStudy (phi):

for 74 b), there are 3! permutations 3 ways to pick the first color, 2 for the 2nd and 1 for the 3rd: 3*2*1 or 3!

OpenStudy (phi):

why do you think a is 3C2 ?

OpenStudy (anonymous):

b) \[\frac{ 3! }{ (3-2)! }\]

OpenStudy (anonymous):

a) the question ask how many combination than is C

OpenStudy (anonymous):

b) the question ask how many permutations than is P

OpenStudy (phi):

yes, but in a) it is 3C3 not 3C2

OpenStudy (phi):

b) is 3P3 which is 3!/0! = 3! = 6

OpenStudy (phi):

or for b) you can count how many there are by looking at the figure. It shows 6 permutations

OpenStudy (anonymous):

FIRST draw is 3 color , SECOND DRAW IS 6 , each color out come 2 3 C2 ?

OpenStudy (phi):

the bag has 3 marbles. you choose 3 of them so it is 3 choose 3 or 3C3

OpenStudy (anonymous):

thanks , can you hep 75?

OpenStudy (phi):

for 75 a) do you know the total number of permutations ?

OpenStudy (anonymous):

\[\frac{ 10! }{(10-5)! }?\]

OpenStudy (phi):

10! permutations total (we don't care about intermission or how many are before the intermission) next, we put your rival and you in position 4, 5 that leaves 8 people to go in the other positions, or 8! ways the answer is 8!/10! = 1/(10*9) = 1/90

OpenStudy (anonymous):

how you get 8?

OpenStudy (phi):

you are told you and your rival are in fixed positions. we have to count all the ways the other 8 people are arranged. (that would be 8!)

OpenStudy (anonymous):

you are the last and you are rival performs immediately before you is mean , last and before is 2 ? than 10 -2 = 8?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thanks , b) you are not the first than 10-1=9 9!/10! ?

OpenStudy (phi):

9!/10! is the number of ways for you to be first. (we fix you in position 1, and find 9! ways to arrange the other 9 people) to get the number where you are *not* first, is do 1 - 9!/10!

OpenStudy (phi):

you should get 1 -1/10 = 0.9 for b)

OpenStudy (anonymous):

don't understand part b

OpenStudy (anonymous):

how you subtract 1 ?

OpenStudy (phi):

if there are 10 people, and we put you first that leaves 9 people that we can arrange. so 9! ways for you to be first. ok with that?

OpenStudy (anonymous):

the question as you are not first , why you said you first ?

OpenStudy (phi):

the probability for being first is easy to figure out and we know Pr(being first) + Pr(not first) = 1 (one or the other must happen) so Pr(not first) = 1 - Pr(first)

OpenStudy (phi):

if you want to figure out Pr(not first) we could do it this way: out of 10 people, 9 can be in the first position (we have to leave you out) there are now 9 people and 9 positions left, and 9! ways to fill them so 9 * 9! divide by 10! (number of ways to arrange 10 people) we get \[ \frac{9 \cdot 9!}{10!} = \frac{9}{10} = 0.9 \]

OpenStudy (anonymous):

you are great, are you Math teacher ?

OpenStudy (phi):

no

OpenStudy (anonymous):

Can you tutor me sometime I will Pay ?

OpenStudy (phi):

you don't have to pay. but I will help when I am around.

OpenStudy (anonymous):

no, I love to pay I not compotable ask help but don't pay, I can pay a little bet to OK,can I email ?

OpenStudy (anonymous):

I will send you a check